### v1

# BIV-Priv-Seg: Locating Private Content in Images Taken by People With Visual Impairments

Abstract
Individuals who are blind or have low vision (BLV) are at a heightened risk of sharing private information if they share photographs they have taken. To facilitate developing technologies that can help preserve privacy, we introduce BIV-Priv-Seg, the first localization dataset originating from people with visual impairments that shows private content. It contains 1,028 images with segmentation annotations for 16 private object categories. We first characterize BIV-Priv-Seg and then evaluate modern models’ performance for locating private content in the dataset. We find modern models struggle most with locating private objects that are not salient, small, and lack text as well as recognizing when private content is absent from an image. We facilitate future extensions by sharing our new dataset with the evaluation server at https://anonymous.com.
1 Introduction
People who are blind or have low vision (BLV) regularly share photographs they have taken. They do so for multiple reasons including to stay connected on social media [1, 5, 48] and to receive assistance with everyday visual tasks [13, 2], such as recognizing objects and reading mail.
A concern is that many images from BLV photographers contain private content. Underscoring this concern, one study showed that privacy leaks occurred for over 10% of more than 40,000 images shared by BLV photographers using a visual assistance technology (i.e., the VizWiz application) [18]. Such privacy leaks are concerning regardless of whether they are seen by remote humans providing visual assistance or retained by companies [45, 46], who may subsequently use such data for training AI models and for accruing licensing revenue from other end users.
Figure 1: We introduce BIV-Priv-Seg, a dataset containing segmented private objects. An example is shown from each category.
To facilitate the development of privacy-preserving technologies, we introduce the first localization dataset that originates from people with visual impairments and shows private content. Called BIV-Priv-FewShot, it extends prior work’s [44] collection of images that were taken by BLV photographers of (non-personal) private content. For each of the 1,028 images in the dataset, we collected human-annotated segmentations for every instance of the 16 photographed private object categories. This culminated in $967$ instance segmentations. Examples of annotated images from every private object category are shown in Figure 1.
We also characterize BIV-Priv-Seg and compare it with five existing few-shot localization datasets to reveal what makes it similar to prior work as well as unique. We found that differentiating aspects of BIV-Priv-Seg include no target objects in 19% of the images, objects exhibiting greater variability in terms of how much of an image they occupy, objects residing more often on image borders, and objects more frequently containing text.
Finally, we benchmark three modern few-shot localization models and vision-language models to reveal their current strengths as well as where improvements are needed. Our fine-grained analysis reveals they struggle most at locating objects that are not salient, small, and lack text as well as recognizing when private content is absent from an image. We expect this work will facilitate the development of models that can handle a greater diversity of real-world challenges and also benefit a larger audience encountering similar privacy challenges, such as children and individuals with cognitive impairments.
2 Related Work
Computer Vision Datasets.
A critical component for successfully developing computer vision models is the availability of human-annotated datasets to support reproducible, quantitative evaluation. A challenge for developing privacy-aware technologies is that most datasets lack private content due to intentional filtering efforts, including for popular, mainstream datasets such as VOC [12], ImageNet/ILSVRC [41], COCO [29], LSUN [51], and Open Images [27]. Only relatively small collections of private images (e.g., thousands of examples) have been successfully curated for public use, including by scraping them from photo-sharing websites (i.e., Flickr, Twitter) [36, 37, 53] and hiring BLV photographers to photograph (non-personal) private content [44]. Extending the latter prior work, we introduce the first publicly-available dataset originating from BLV photographers that shows segmented private content. This effort extends the broader research community’s inclusivity efforts to increase the availability of datasets based on images taken by people with vision impairments [3, 6, 8, 9, 19, 20, 25, 31, 17, 18, 52].
Given that it is non-trivial to curate a large-scale collection of private images for supporting the development of privacy-aware models, since the very nature of such data means it typically cannot be saved or shared, we focus on introducing a dataset challenge that facilitates the development of zero-shot and few-shot models (and so requires a small amount of training data). Accordingly, our dataset challenge will complement prior work that proposed few-shot object detection and instance segmentation challenges, including VizWiz-FewShot [47], PASCAL-$5^{i}$ [43], COCO-$20^{i}$ [35, 32], and LVIS [34], and FSOD [14]. Our dataset offers unique challenges compared to existing datasets, including images lacking the target object, a wider variability of object sizes, a higher prevalence of objects residing on image borders, and a higher prevalence of objects with text.
Zero-Shot and Few-Shot Object Localization Models.
Zero-shot and few-shot methods that can locate objects in images span more traditional few-shot object localization models and more recent vision-language models. Traditional methods include those designed to only achieve strong performance on novel object categories [38, 32, 10, 50, 28, 30, 21, 11, 15, 33, 34] as well as those designed to also maintain performance on base object categories [4, 16] (i.e., generalized/incremental setting). A new breed of relevant models also emerged in the past year: large multimodal models (LMMs) that can generate a textual description alongside a visual grounding (i.e., segmentation mask) specifying where the described content is located [39, 49]. [^1] Also related is a framework introduced in 2024 [42] for adapting vision-language models (VLMs) to identify privacy levels of conversational scenarios. However, their work does not support locating private content, restricting its scope for potential privacy-preserving technologies. LMMs offer the advantage of being more interactive, making them particularly appealing for assistive tools for the BLV community. We benchmark a total of three models (2 traditional and 1 LMM) to assess how well modern methods perform on our new dataset. While we observe the best performance from LMMs, all models still fall below human performance. Our results suggest that promising future directions for research are to improve models’ abilities to handle objects that are not salient, small, and lack text as well as to recognize when private content is absent from an image.
3 BIV-Priv-Seg Dataset
We now introduce our dataset for locating private objects in images taken by people with vision impairments.
3.1 Dataset Creation
Data Source.
We leverage the BIV-Priv dataset [44], which consists of 1,176 images taken by 26 BLV photographers of 16 types of private objects. All private objects were ‘props’, meaning that the objects lacked any private information pertinent to the photographer (e.g., a fake medical bill rather than a real one). Each BLV individual took pictures with each object type positioned to appear in the image’s foreground as well as its background. We leverage all images, including those that lacked the target private content, to reflect the unique challenges faced by BLV photographers who often do not successfully capture the content of interest in their images
Category Selection.
We chose $16$ private object categories, after slightly modifying the original categories proposed in BIV-Priv [44] (as discussed in the Supplementary Materials). Our final set of categories is as follows: “bank statement”, “bill or receipt”, “business card”, “condom box”, “condom packet”, “credit or debit card”, “doctor’s prescription”, “pill bottle”, “letter with address”, “local newspaper”, “medical record document”, “mortgage or investment report”, “pregnancy test”, “pregnancy test box”, “tattoo sleeve”, and “transcript”.
Data Filtering.
We filtered all images containing any personally identifiable information (PII) from the photographer in a multi-step approach. After receiving approval from Institutional Review Boards (IRB), two authors who were trained to recognize all ‘prop’ private objects and the definition of PII reviewed every image independently. When both individuals flagged an image as containing PII, it was filtered. When only one flagged an image, it was discussed by three authors to make a final decision. This process yielded $1,028$ images suitable for public release.
Annotation Task.
We leveraged the instructions and annotation protocol established by prior work (i.e., VizWiz-FewShot [47]) to enable fair comparison with it. To create instance segmentations, annotators clicked a series of points to generate polygons, dragged a polygon’s vertices to refine it, and created holes by drawing ‘interior’ polygons in ‘exterior’ polygons. We leveraged the annotation platform, Supervise.Ly [^2] https://supervisely.com/ to collect all segmentation annotations.
Annotation Collection.
We next localized all instances of the 16 private object categories in every image. To collect high-quality annotations, we relied on three in-house annotators who were authors of this work and so intimately familiar with the images. Each individual had a 1-hour 1-on-1 training session to learn about the annotation interface and then passed a test demonstrating they could accurately annotate a sample image with the annotation tool. Each image was annotated by two annotators and we used $IoU$ scores to determine how to establish a ground truth segmentation per image. When $IoU\geq 0.8$, we randomly chose one annotation for the ground truth. Otherwise, the three authors reviewed the pair of annotations to choose one as the ground truth (or, in exceptional cases, discarded both annotations). [^3] Only $15\%$ of the images required manual review, with $93\%$ having one of the annotations deemed correct. This result culminated in 967 instance segmentations.
Figure 2: Examples from our BIV-Priv-Seg dataset illustrating its unique aspects, specifically a high variability for the proportion of images occupied by the objects (with many objects occupying much of their images’ borders), high prevalence of text in objects, and inclusion of images lacking target objects.
3.2 Dataset Analysis
We now characterize BIV-Priv-Seg and how it compares to the following mainstream few-shot localization datasets: VizWiz-FewShot [47], PASCAL-5i [43], COCO-20i [35], LVIS [34] and FSOD [14]. Our findings will show that our dataset offers several unique aspects compared to existing datasets, which are exemplified in Figure 2.
Global Dataset Statistics.
We analyze three properties for each few-shot localization dataset: the number of annotated objects per image, the number of disconnected areas an object contains (i.e., segments), and the percentage of objects containing text [^4] Objects are flagged if Microsoft Azure’s Optical Character Recognition (OCR) API detects text within the image, after masking out all content except for the instance segmentation.. PASCAL-5i and FSOD are excluded from the latter two metrics since their annotations only contain bounding boxes and so lack the necessary segmentation information to count segments and locate texts. Results are visualized in Figure 3.
Figure 3: Comparison of our dataset to four existing few-shot localization datasets with respect to the (a) number of annotated objects per image, (b) number of segments (i.e., disconnected areas) per annotated object, and (c) percentage of annotated objects containing text.
Figure 4: Comparison of our dataset to four existing few-shot localization datasets with respect to the (a) border rate, (b) center deviation, (c) image coverage, and (d) boundary complexity. PASCAL-5i and FSOD are excluded for boundary complexity as they lack the necessary segmentation annotations to calculate this metric.
When observing the histogram showing the percentage of images containing different numbers of annotated objects (Figure 3(a)), we observe two key distinctions of our dataset. First, our dataset is the only one with images lacking objects of interest, with this occurring for 18% of the images. These reflect both when no target object is visible and when the target object is a paper positioned on its blank side, rather than the side showing the document’s text (and so private text as well). This amount is similar to that reported by prior work [9], where roughly 20% of BLV photographers’ images lacked the visual evidence needed to answer their visual question despite showing recognizable content [20]. We retain such images to reflect that the photographers intended to capture an object of interest in such images [44], highlighting the challenge of dealing with target-less samples. This is to encourage modern algorithms to achieve a good true negative rate, which is critical for many real-world applications. To support fair dataset comparisons, we exclude these images when analyzing our dataset for the rest of this section. The second key distinction of our dataset is that it rarely has more than one annotated object per image. We attribute this to the distinct photography goal for BIV-Priv [44], where the goal for every image was to photograph a single object, whereas other datasets such as COCO-20i were intentionally curated to show complex scenes showing more than one object (i.e., none-iconic images). When observing the histogram showing the percentage of images containing different numbers segments (i.e., disconnected areas) (Figure 3(b)), we observe that our BIV-Priv-Seg dataset typically exhibits only one segment per object. VizWiz-FewShot, in contrast, more often has the most with up to $122$ segments in an object. While an object itself exists as a unified entity in the real world, multiple counts of segments arise in an object under two circumstances: (1) whenocclusions break objects into more than one region and (2) when objects are partially outside the frame of the image so there are disconnected regions of an object. Our dataset limits these scenarios, thereby permitting more focused model benchmarking analysis for other challenges that can arise, as discussed in this section.
When analyzing the prevalence of text in objects, we find it is the highest for our dataset, occurring for $74.5\%$ of annotated object instances (Figure 3(c)). We attribute this greater prevalence of text to the intrinsic nature of the private categories, where textual information often is the private content. For instance, paper-based categories such as “bill or receipt” and “doctor’s prescriptions” can show names, addresses, and more. Consequently, our dataset provides an opportunity to shift algorithm developers’ focus to more heavily prioritize analyzingtext on objects.
Localized Instance-Based Statistics.
We next analyze the properties of annotated object instances. We apply metrics to each dataset’s annotations as is, whether they are segmentations or bounding boxes. For every instance, we compute the following four metrics:
  1. 1.
    Border rate: ratio of the number of pixels in the object boundary touching the image border to the total number of pixels on the image border. An object touching the image border likely signifies that part of the object is outside the frame.
  2. 2.
    Center deviation: ratio of the distance between the image center and the object center to the distance of the image center to its furthest pixel in the image. This indicates how far an object center’s from the image center and so the traditional photographer’s bias of centering visual content in the field of view.
  3. 3.
    Image coverage: ratio of the object area to the image area. This reveals the size of an object in the frame.
  4. 4.
    Boundary complexity: one minus the isoperimetric inequality, which is the ratio of the area of an instance to the length of its perimeter. A lower isoperimetric inequality value indicates a more complex boundary. This metric is informative only for segmentation annotations, so we exclude datasets with only bounding box annotations.
Results are shown in Figure 4, with our dataset exhibiting a couple of important differences. First, our dataset tends to have the largest border rate, followed by VizWiz-FewShot (Figure 4(a)). This finding aligns with prior work’s quality assessments of BLV photographers’ images, where inadequate framing including from out-of-frame areas was found in more than half of the (VizWiz) images [9]. Second, our dataset tends to exhibit objects with the most image coverage (Figure 4(b)), thereby encouraging the development of algorithms that can handle an even wider range of object sizes. It is worth noting, that the positive correlation between border rate and image coverage likely arises for the same reason: closer proximity to the target object results in a larger object size within the image and so increased likelihood for the object to be partially out of frame.
4 Algorithm Benchmarking
We now analyze modern methods’ performance on our new dataset. We conduct analyses for both traditional few-shot localization algorithms as well as vision-language models (VLMs) in the zero-shot setting.
4.1 Few-shot Localization Algorithms
We benchmarked the top-performing few-shot object detection (FSOD) and few-shot instance segmentation (FSIS) models for which publicly available code could be successfully deployed on modern GPUs: DeFRCN [38] for FSOD and YOLACT [7] for both FSOD and FSIS. [^5] We adopted the hyperparameters and training pipelines from the prior work [38] for DeFRCN and the prior work [47] for YOLACT. We trained the models for base categories using the larger-scale VizWiz-FewShot [47], since those examples also originate from BLV individuals, and then fine-tuned the models on the categories in BIV-Priv-Seg.
Training and Evaluation Splits.
The experiments were conducted using 4-fold cross-validation, where the 100 categories in VizWiz-FewShot were divided into four sets, with the combination of three sets (75 categories) serving as base categories in each fold. This setting matches prior work [47] for a fair comparison to VizWiz-FewShot.
Evaluation Metric.
We evaluated the models using the mean Average Precision (mAP), which is the average precision across a range of intersection over union (IoU) thresholds from 0.5 to 0.95 in steps of 0.05. This metric aligns with prior work [47, 38, 33, 22]. We also present results with respect to $mAP_{50}$, where only threshold $0.5$ is used since this approach facilitates the comparison with datasets such as Pascal VOC.
Baseline Datasets.
For comparison, we also analyzed each trained model’s performance for other few-shot localization datasets: PASCAL-$5^{i}$[43], COCO-20i[35, 32], and VizWiz-FewShot [47].
Overall Performance.
Results are in Tables 1666The benchmarking results reported for DeFRCN performance on PASCAL-$5^{i}$ is presented in the format specified in [43] with three data splits. To compare these results to the 4-fold cross-validation format, we average the scores of the three pre-defined splits. and 2.
Overall, we observe similar DeFRCN performance on our dataset to that observed for the baseline datasets (Table 1). With respect to the mAP scores, our dataset appears to be easier compared to VizWiz-FewShot and COCO-20i, while it is harder thanPASCAL-5i when considering the AP50 scores. A commonality across all the datasets is that the DeFRCN’s performance still is quite poor, falling short of what might be practically useful in an end-user-facing tool. Overall, these findings underscore the potential benefit of our dataset in helping advance modern few-shot learning algorithms.
Metric Dataset   \   $k=$ 1 3 5 10
BIV-Priv-Seg (ours) 12.7 19.4 23.8 29.7
mAP VizWiz-FewShot [47] 4.0 7.4 10.5 12.6
COCO-20i [35] 9.3 14.8 16.1 18.5
BIV-Priv-Seg (ours) 21.7 30.3 36.4 44.7
AP50 VizWiz-FewShot [47] 6.9 12.8 17.6 23.5
PASCAL-5i [43] 44.0 53.6 57.4 55.4
Table 1: Results from DeFRCN [38] on our dataset and other few-shot instance segmentation datasets with respect to two evaluation metrics: mAP and AP50.
Overall, we observe YOLACT performs poorly on all datasets with slightly better results on our dataset than that observed for the baseline datasets (Table 2). For instance, the AP50 scores of the segmentations of our dataset are 24.46 and 38.53 for $k={1,5}$ compared to 19.0 and 21.7 of COCO-20i and 12.03 and 24.99 of VizWiz-FewShot. We suspect this finding stems from our private objects having a notably higher prevalence of text and image coverage, as prior work [47] has suggested such trends.
FSOD mAP FSOD AP50 FSIS mAP FSIS AP50
Dataset   \   $k=$ 1 3 5 10 1 3 5 10 1 3 5 10 1 3 5 10
BIV-Priv-Seg (ours) 18.1 23.3 28.7 40.0 24.7 31.7 39.4 52.4 20.0 25.6 31.1 44.4 24.5 31.3 38.5 52.2
VizWiz-FewShot [47] 10.7 11.7 22.6 32.5 13.2 14.9 26.9 43.2 10.0 10.5 20.3 28.7 12.0 12.8 25.5 36.4
COCO-20i [35] - - - - 19.0 - 21.7 - - - - - 17.1 - 18.9 -
Table 2: Results from YOLACT [7] on our dataset and other few-shot instance segmentation datasets with respect to mAP and AP50. Results are presented for both instance segmentation and object detection.
Figure 5: Qualitative results from DeFRCN in the 1-shot setting for object detection (bounding boxes), YOLACT in the 1-shot setting for instance segmentation (masks) and object detection (bounding boxes), and GLaMM using the prompt I scenario for the zero-shot setting. As a point of reference, ground truth results are also provided.
Qualitative results from the 1-shot settings are shown for both models in Figure 5. In these examples, we can see that YOLACT and DeFRCN tend to miss the same objects, while YOLACT is able to locate some objects that DeFRCN misses.
Fine-Grained Analysis.
We next perform fine-grained analysis of the models by dividing the test examples into subsets based on the following characteristics also employed by prior work [47]: object size and presence of text. We categorize objects into small, medium, and large sizes, by sorting all instances in BIV-Priv-Seg based on their image coverage and then equally divide them into three groups. To detect the presence of text, we used OCR on background-masked instance images to flag whether text is present. Results are provided in Table 3 with respect to mAP scores. The trends of AP50 scores are consistent with that of the mAP scores and so are excluded from the table.
For object size, our findings for YOLACT reinforce those of prior work [23, 54, 47], showing that it performs worse for smaller objects. However, DeFRCN’s results indicate that the models tend to perform better on medium-sized objects. The observation of these differences in model trends provides an insight for future BLV applications that the unique object size characteristics could play a crucial role in the design of these models.
With respect to presence of text, we observed that both models consistently perform better on objects with text than without text. This finding supports the observations in previous work [47] that texts could serve as strong features for the models to localize objects in few-shot learning. This emphasizes the advantages of localization of private objects based on the high prevalence of text, highlighting a strong foundation for privacy applications in the area.
Object Size Text
Model $k$ S. M. L. Yes No
$1$ 12.41 22.60 26.53 24.56 11.24
YOLACT $3$ 18.18 31.54 36.89 34.75 12.86
(FSIS) $5$ 22.90 36.97 44.39 42.32 18.31
$10$ 29.50 55.19 63.28 58.56 31.98
$1$ 8.70 15.50 13.33 15.63 4.95
DeFRCN $3$ 11.71 23.21 23.29 23.76 8.08
(FSOD) $5$ 15.73 26.49 25.11 28.13 11.45
$10$ 17.89 34.67 30.32 34.58 15.02
Table 3: Fine-grained analysis of the performance of FSIS and FSOD models on BIV-Priv-Seg presented in mAP.
4.2 VLM Localization
We next benchmarked the top-performing, off-the-shelf vision language model that supports visual grounding in a zero-shot settings: GLaMM [39]. We explored prompting the model in the following three ways:
Prompt I. Privacy-agnostic(category specific): Each category’s name is inserted into the following prompt: “Segment it if there is a {category name} in the given image, otherwise say there is no {category name}.” This scenario is very general, rather than specializedin the concept of privacy.
Prompt II. Privacy-aware: We use the following generic prompt “Segment it if there is a private object in the given image, otherwise say there is no private object.”, where the specific category name is replaced with the generic phrase “private object”. This prompt captures is intended to capture model’s ability to understand the concept of privacy.
Prompt III. Privacy-explained: To provide the context of privacy, we provide the list of private categories in our dataset and then request the segmentation as follows: “Private objects include {all category names}. Segment it if there is a private object in the given image, otherwise say there is no private object.” We intentionallydesigned these model prompts toallow for a negative response regarding the presence of the target, meaning that the model does not return a segmentation result. We achieved this by extending the grounding prompts from previous work [39] from“Segment something” to “Segment it if something, otherwise say there is no something”.
Evaluation Method.
We introduce a two-step evaluation method for handling when target objects are absent from an image. First, we categorize the images into a binary classification task with two classes based on whether private objects are found in the images. Then, we evaluate the localization results on the target-containing predictions.
Privacy classification: We first evaluate the model performance on two levels: (1) Semantic level prediction, which is based on the language used in the conversation, where positive responses such as “yes, here is the mask” or “sure, here is the mask” indicate the presence of a target, and negative responses such as “no, there is no…” indicate the absence of a target; and (2) Mask level prediction, which is determined by whether any pixel in the generated mask has a positive prediction. The results are presented in accuracy, recall, and true negative rate. Accuracy shows the overall performance; recall reflects the percentage of correct predictions among the target-containing images; true negative rate reflects the percentage of correct predictions among the target-less images. Segmentation: For the images that are classified as private on both semantic and mask levels, we then perform localization. The mIoU and cIoU metrics are used to evaluate the semantic segmentation results across all categories in each scenario. As most of the images in this dataset only contain one target object, we convert those segmentation results into instance segmentation format and evaluate the models in AP50 and Recall metrics. This two-step evaluation highlights the model’s ability to identify whether there are private objects in the image and its localization ability.
Overall Performance.
Below, we present the results of classification and segmentation, respectively.
The results of both the semantic level and mask level privacy classifications are shown in Table 4. Two key findings emerge. First, on the semantic level, the model consistently generates positive responses across all scenarios, producing masks for target objects in every given image. However, some of these generated masks contain negative results, indicating that while the model confirms the presence of a private object at the conversational level, the masks do not actually predict any pixels corresponding to the private object. This inconsistency suggests that the model lacks a robust text-mask alignment mechanism. Second, despite some negative predictions at the mask level for prompt II and prompt III, the true negative (TN) rates are still extremely low. This indicates the model’s inability to accurately identify when there are no target objects in the image. Prompt I results, which are aggregated across all object categories, show that at least one category is predicted to contain the target in every image in BIV-Priv-Seg.We suspect that bothissues stem from biases in the training data, where there is a lack of target-less samples.
Semantic level Mask level
Prompt Acc R TN Acc R TN
Prompt I 81.32 100.00 0.00 81.32 100.00 0.00
Prompt II 69.26 83.97 5.21 81.32 100.00 0.00
Prompt III 63.71 74.28 17.71 81.32 100.00 0.00
Table 4: Classification performance of GLaMM on BIV-Priv-Seg in classifying the images into target-containing and target-less groups on both the level of semanticsand masks. Results are presented in accuracy (Acc), recall (R), and true negative rate (TN) as percentages.
Results for the target-containing group are shown in Table 5, withqualitative results in Figure 5. We first compared the performance of GLaMM on our dataset to those on VizWiz-FewShot [47], GranD [39], refCOCO, refCOCO+, and refCOCOg [24]. The results show that the performance of our dataset is comparable to existing datasets. Specifically, the performance on refCOCO, refCOCO+, and refCOCOg is better than BIV-Priv-Seg (prompt I), while the performance on GranD and VizWiz-FewShot is worse. This indicates that state-of-the-art VLMs have the potential to effectively ground private objects in images, achieving performance comparable to other established datasets. We then compared the performance of prompts I, II, and III to explore the model’s ability to comprehend privacy. We observed that prompt I achieved the highest performance, indicating that when given a specific object category, the model has the potential to locate private objects accurately. Conversely, prompt II had the worst performance. Even though the categories in BIV-Priv-Seg are perceived as private by humans and the semantic meaning of prompt II encompasses these categories, the model struggled to localize the private objects as effectively as it did with prompt I. This suggests that the model’s understanding of privacy is less nuanced than human perception. Interestingly, while prompt III attempted to mitigate this semantic ambiguity by providing an explanation of privacy, it still performed poorly compared to prompt I. Despite the initial definition of privacy in prompt III, the model failed to fully comprehend and execute the instructions. This indicates that the model is highly sensitive to the sentence structure of prompts, which may reflect biases in its training data.
Dataset mIoU cIoU AP50 Recall
BIV-Priv-Seg (ours, prompt I) 68.0 51.0 44.7 55.9
BIV-Priv-Seg (ours, prompt II) 11.3 -26.1 3.1 7.8
BIV-Priv-Seg (ours, prompt III) 20.6 -0.5 9.4 14.8
VizWiz-FS [47] (prompt I) 15.6 -16.1 2.2 5.6
GranD (validation) [39] 66.3 - 30.8 41.8
GranD (validation) [39] 65.6 - 29.2 40.8
refCOCO (testA) [24] - 83.2 - -
refCOCO (testB) [24] - 76.9 - -
refCOCO+ (testA) [24] - 78.7 - -
refCOCO+ (testB) [24] - 64.6 - -
refCOCOg (test) [24] - 74.9 - -
Table 5: Overall performance of GLaMM on the target-containing group of BIV-Priv-Seg compared to existing grounding datasets. BIV-Priv-Seg is evaluated in three prompt types, and VizWiz-FewShot is evaluated in the privacy-agnostic scenario only as the dataset contains non-private categories only. The results of GranD, refCOCO, refCOCO+, and refCOCOg are adopted from previous work [39].
Fine-Grained Analysis.
We conducted analyses similar to object size and presence of text as we did for few-shot localization algorithms and additional analysis on object saliency. Becausee observed from qualitative results that VLMs tend to segment the salient object in the images, we also evaluated based on whether the target object is the salient object. To do this, we identified whether a target object is salient by applying a salient object detection model [26] that has shown to yield comparable performance to humans for images taken by people with visual impairments (i.e., for the VizWiz dataset) [40] and then applyinga $0.5$ threshol to its intersection over union (IoU) with the target object. This means that when over half of the masks are overlapped, the object is considered salient in our analysis. The results of object size and presence of text align with that of few-shot localization, and we found in the results of object saliency that the model performs better on salient objects.
Saliency Object Size Text
metric Yes No S. M. L. Yes No
mIoU 85.1 57.1 37.8 79.2 85.2 73.8 51.2
cIoU 75.7 35.3 23.7 67.1 80.4 61.2 21.4
AP50 79.8 37.9 19.5 73.9 84.1 60.5 30.3
Recall 79.9 44.5 24.9 74.0 82.8 64.4 38.2
Table 6: Fine-grained analysis of the performance of GLaMM on BIV-Priv-Seg in prompt I scenarios.
Results are presented in Table 6. For object size, we see a consistent trend of performance growth from smaller to larger size; for presence of text, the performance on objects with texts is consistently higher than those on objects without texts. These results correspond to previous findings and reflect the existing gaps between modern algorithms and applications. Concerning object saliency, we observe a strong correlation between saliency and segmentation accuracy. The localization performance is consistently and significantly better on salient objects than on non-salient objects. This relation implies that the modern VLMs rely on an object’ssaliencyfor grounding, as opposed to a complete reliance on the features of the target objects, while also suggesting a potential bias in the existing training data.
5 Conclusions
While few-shot and zero-shot localization algorithms have made significant strides in recent years, their application to privacy preservation has been limited due to the scarcity of publicly available, annotated images containing private visual information. Yet, such algorithms bring significant potential for building privacy-aware assistive tools because they empower each user or developer to relatively easily embrace distinct definitions of what is private.
To fill this gap, we introduced a new few-shot localization dataset challenge that contributes to the development of privacy-preserving technologies for people who are blind or with low vision: BIV-Priv-Seg. Our extensive analysis of modern algorithms on BIV-Priv-Seg reveals they struggle most with private objects that are small, lack text, and are non-salient. Despite the current gaps between models’ capabilities and real-world privacy preservation needs, the performance of the benchmarked models provides promising prospects for promoting a safer user environment. We offer our work as a valuable foundation for developing more flexible and personalizable privacy-aware solutions that can cater to individual needs.
We envision two important directions for future work. First, is supporting personalized identification of private objects on one’s own devices, where users train models to recognize their private objects in any new images or videos on-device. A critical direction to make this idea a reality is developing such models to be compact so that they can run on resource-constrained devices. A second promising research direction is infusing privacy-aware capabilities into modern VLMs since BLV users already rely on such conversational agents (e.g., Be My AI [13], Microsoft’s Seeing AI [2]). This will enable them to seamlessly realize privacy-preserving capabilities without considerably changing their workflows. Potential interactions include identifying private content in images, obfuscating the information, and providing detailed privacy status updates through a conversational interface. Critical research directions to bring this idea to reality include reconciling VLMs’ inconsistencies between text and mask predictions while also infusing them with the ability to recognize when requested objects are absent from an image.